SPECTROSCOPY OF CARBON CHAINS AND THEIR IONS

\underline{J. P. MAIER}, Institute for Physical Chemistry, University of Basel, Klingelbergstrasse 80, CH-4056 Basel, Switzerland.

The electronic spectra of neutral carbon chains, their anions and cations are being observed in the gas-phase. Three different approaches are used. The transitions of the chain radicals, C_{n}H n=6,8,10 have been detected by cavity ring down spectroscopy in a slit jet discharge. Carbon chain anions of the type C_{n}^{−}, C_{n}H^{−} have been studied by a two colour resonant photodetachment approach. The electronic transitions of the polyacetylene cations, like HC_{6}H^{+} and its cyanoderivative HC_{4}CN^{+}, have been measured at high resolution in cell and jet discharges using frequency modulation absorption spectroscopy.